Reports 1-1 of 1 Clear search Modify search
Detector Characterisation (Spectral lines)
direnzo - 10:45 Thursday 20 March 2025 (66411) Print this report
Excess Noise at Violin Mode Frequencies After March 18 Maintenance

Since the maintenance on March 18, an excess of noise has been observed at the frequencies of the violin modes (430-450 Hz, and harmonics).

Figure 1 shows the spectrogram from the past three days in the 400-500 Hz range, clearly highlighting the increase in noise since March 18.

Figure 2 presents the ASD at three different time intervals.

Looking at the results of the daily BruCO runs, it seems there is a very high coherence with the channels V1:SDB2_POWERSUPPLY_DBOX_LEFT_DOWN_*. Compare the results before the maintenance (Mar 17), and after it (Mar 19).

Images attached to this report
Comments to this report:
mwas - 13:23 Friday 21 March 2025 (66421) Print this report

SDB2_POWERSUPPLY_DBOX_LEFT_DOWN_* are channels for the power supply on which the main B1 photodiode are plugged. The large lins at ~440Hz are of known physical origin, these are the violin modes of the mirror suspension fibers, and these have been excited by work on the north end suspension during the maintenance.

The fact that there is coherence between h(t) and SDB2_POWERSUPPLY_DBOX_LEFT_DOWN_* means that these power supply channels are not safe for large amplitude signal in B1. A signal on the light, is causing a signal on the power supply, as the amount of current that the photodiode is drawing from the power supply fluctuates with the signal.

salvador - 17:06 Friday 21 March 2025 (66424) Print this report
Search Help
×

Warning

×